# TD-Orch: Scalable Load-Balancing for Distributed Systems with Applications to Graph Processing

### v1

@@ 2.2. Prior Scheduling Strategies
Direct Push. Another common strategy in distributed settings is to offload each task to the machine that stores its required data chunks. However, machines holding “hot” data chunks incur high communication cost (for receiving the tasks) and high computation work (for executing the tasks it receives), resulting in severe load imbalance. Moreover, the $O(1)$ data chunks required by a task may be spread across multiple machines, in which case there is no one machine holding all its required data chunks.
@@
Sorting-based. Prior work on the Massively Parallel Computing (MPC) model (im2023massively) has proposed theoretically-efficient solutions to the orchestration problem (e.g., (im2023massively; goodrich2011sorting)). The procedure begins by sorting tasks according to the addresses of their required data chunks. Then, MPC broadcasting (ghaffari2019massively) is used to distribute each data chunk to the corresponding tasks. After task execution, a reverse broadcasting step writes the results back to the data chunks, followed by a reverse sorting step that restores tasks to their original order. It has been shown that the computation work is load balanced, and the communication cost per machine is $\Theta(\frac{n}{p}\log_{\frac{n}{P}}{P})$ (fish2015computational), which is optimal (roughgarden2018suffles). (Note that sorting is used here solely for grouping tasks by their requested data, and hence semi-sorting would suffice.)
Although these MPC schemes are asymptotically optimal, we will show below (Section 2.5) that the constant factors are non-negligible in practice.
2.3. TD-Orch
To address the challenges of direct push or pull, we must handle potential contention. Our method, TD-Orch, exploits the two directions of data flow: push, where (the context of) a task is sent to the machine(s) storing its required data chunks, and pull, where data chunks are fetched to the machine where the task resides. Such bidirectional strategies (referred to as distributed push-pull) are rarely emphasized in prior work. Only a few prior works demonstrate that this strategy can significantly improve load balance, but they are limited to PIM settings (kang2022pimtreevldb; kang2023pimtrie; zhao2025optimal) and rely on an extremely powerful centralized coordinator (the host CPU). Our goal is to develop a more widely applicable push-pull technique that does not require a centralized coordinator.
Ideally, when contention information is available on any machine on the fly, distributed push-pull merges redundant information on one side (task/data) and dynamically chooses the flow direction toward the contended end. This dynamic decision should balance load with low communication.
Orchestration Algorithm. TD-Orch unifies information collection and push within a single substage, where the task locations for pull are also decided. Bounded-size messages are enabled for dynamic decisions, even when machines lack direct visibility into remote states. Its main algorithm to enable load-balanced task-data orchestration consists of the following four substages in each stage:
@@ 2.3. TD-Orch
    Contention Detection: The information of each task is sent to the machine storing the target data, along a communication forest (§2.3.1) from leaves to root. Contention information on each data chunk is collected during this substage, in the form of a meta-task (§2.3.2).
@@
    Task-Data Co-location: A meta-task inherently provides a highly-compressed tree for communication. Data copies are broadcast along this tree from root to leaves. (See §2.3.3.)
@@
    Data Write-backs: The updated value for any data modified by an executing task is written back to the original data location. (See §2.3.4.)
2.3.1. Communication Forest
During contention detection, load balance must be preserved even before the system identifies contention. This requires that each machine only sends or receives bounded-size information, while avoiding excessive communication or round overhead caused by these bounded sizes. To meet these requirements, TD-Orch adopts a tree/forest organization of machines for load-balanced communication, inspired by MPC broadcast trees (ghaffari2019massively).
@@ 2.3.1. Communication Forest
A communication tree rooted at machine $i$ is a balanced $F$-ary tree with $P$ leaves, as illustrated in Figure 2. Each of the $P$ leaf nodes represents a different physical machine, while the internal nodes correspond to virtual transit nodes, each of which is mapped to one of the $P$ machines via a hash function. Notably, a physical machine can simultaneously serve as both a leaf and an internal node. The hash function is globally known to all machines.
Figure 2. An example communication tree with fanout $F=2$ in $P=8$ machines rooted at machine 11. PM stands for physical machines. VM stands for virtual transit machines, whose ID is a tuple of (root ID, BFS ordered ID). Dash lines mean that PMs do not need to send messages to themselves up the trees. In practice, $F\gg 2$.
A communication forest is defined as a set of $P$ communication trees, each rooted at a distinct machine. This structure enables efficient collection of task-related information (such as counting task occurrences) across a batch with near-optimal performance. For each task whose data resides on physical machine $i$, the current machine sends a packed message–containing the data chunk address and the task information–to its parent in the communication tree rooted at $i$. In successive communication rounds, each parent merges the information received from its children and forwards the aggregated result to its own parent in the forest. The stage completes when all information reach their respective roots, requiring $O(\log_{F}{P})$ rounds in total.
2.3.2. Meta-Task
Two additional goals are desirable within the same contention-detection substage. (i) Task contexts without contention should be directly pushed to their data locations, avoiding the need for an extra round. (ii) For data items with contended accesses, the system should also identify where the data must be pulled in the next substage (i.e., the locations of the contending tasks).
To meet these goals, we introduce an auxiliary structure called a meta-task. Meta-tasks are bounded in size to satisfy load-balance requirements. They carry uncontended tasks as they move between machines but exclude contended ones. In addition, their inherent tree structure naturally guides the broadcast routes for data in the subsequent pull substage.
As illustrated in Figure 3, a meta-task is composed of an integer level $l$ (denoted by $L_{l}$) and data contents with constant word-sizes. Each $L_{0}$ meta-task corresponds to a task in the batch and its data content holds all the local context of the task. For $L_{l}$ meta-tasks with $l\geq 1$, the data content is a physical address pointing to the memory address of some machine where a set of $L_{l-1}$ meta-tasks are stored.
@@ 2.3.2. Meta-Task
A meta-task set with chunk size $C$, as shown in Figure 3, is a set structure containing meta-tasks. The number of meta-tasks inside with the same level should not exceed $C$.
@@
A merge operation of two meta-task sets is supported. A new meta-task set with chunk size $C$ is generated, containing all meta-tasks in both of the original sets. If the number of meta-tasks with certain level $l$ in the new set exceeds $C$, all $L_{l}$ meta-tasks will be popped out of the new set and stored locally on the current machine, and then a new $L_{l+1}$ meta-task will be pushed into the set, whose data content contains a physical pointer to where these $L_{l}$ meta-tasks are stored. The merge will execute this process from the lowest level up to highest level, until the number of meta-tasks at each level is at most $C$. An example of merging two sets with chunk size $C=3$ is illustrated in Figure 4.
This merging guarantees that meta-task sets remain bounded in size. Moreover, employing multiple layers of merging ensures that broadcasting in the subsequent pull substage incurs bounded communication per machine, preventing single data from needing to broadcast to numerous children In short, the meta-task hierarchy creates a compressed, communication-efficient broadcast/reduction tree such that (unlike Figure 2) the height and size of the tree adapts to the precise amount of data contention.
Figure 3. A meta-task set with chunk size $C$, with a zoom-in into $L_{0}$ and $L_{l}(l\geq 1)$ meta-tasks contained, and the pointers.
@@
Figure 4. Merging two meta-task sets with chunk size $C=3$. Meta-tasks on lowest level $L_{0}$ are merged first, and then from low-level $L_{1}$ throughout to the highest-level $L_{3}$.
2.3.3. Distributed Push-Pull
Using meta-tasks allow us to send both the counts of tasks that need particular data and implicit information about which set of machines possess such tasks over the communication forest. During the preprocessing, each task is wrapped in an $L_{0}$ meta-task, and meta-task sets are constructed for those tasks accessing the same data chunk. In each communication round the constructed meta-task set is sent in the message in addition to the counts of the tasks. After receiving messages in each round, each parent node merges both the task counts and the meta-task sets. Each machine that stores a data chunk will receive the final meta-task sets and implicitly knows where these tasks are stored.
Since the size of each meta-task set is bounded (though not by constant words), the meta-task collection is still nearly perfectly load balanced. Meanwhile, the hierarchical meta-task sets also provide an implicit tree spanning the subset of machines that possess related tasks, which we denote as the subset tree. When fewer than $C$ tasks target the same data chunk, these tasks are directly pushed to the corresponding machines. If more than $C$ tasks target the chunk, they must access the data copies via a broadcast over the subset tree.
2.3.4. Write Backs
If tasks require write-backs that meet a special merge-able characteristic (defined in 2), then write-backs can be performed efficiently in a bottom-up fashion using the subset trees.
@@ 2.3.4. Write Backs
Discussion. Note that TD-Orch requires only two sweeps over the communication forest, whereas MPC sorting (im2023massively; goodrich2011sorting) incurs more than three passes. This reduction highlights the potential of TD-Orch for strong practical efficiency.
2.4. Formal Guarantees
We use Bulk-Synchronous Parallel (BSP) model (valiant1990bridging; valiant1990general; valiant1988optimally) here to provide theoretical guarantees for our design. A BSP system consists of $P$ machines, each with a sequential processing element (PE) and a local memory without shared global memory. All machines are connected by a point-to-point communication network. The system operates in BSP rounds (supersteps), where periodic barrier synchronizations arer required. The BSP model accounts for the computation time (maximum computation work) and communication time (maximum communication amount) over all machines. Due to using maximum when defining, load balance is critical for achieving good performance. Readers with interests in theory can refer to Appendix A for more details of the model.
@@ 2.4. Formal Guarantees
Our main Theorem 2.1 proves TD-Orch to be asymptotically optimal compared with the lower bounds in (roughgarden2018suffles). Please refer to Appendix B for its proof.
Theorem 2.1 (Main Theorem).
@@
Assume $n=\Omega(\frac{P\log^{3}{P}}{\log\log{P}})$ and $n=O(P^{\epsilon})$ for some $\epsilon>1$. If each of the $P$ machines holds $\Theta(n/P)$ tasks at the beginning of a stage, and write backs are merge-able, then (i) TD-Orch takes $O(\frac{n}{P}\log_{\frac{n}{P}}{P})$ BSP communication time whp  [^1] We use $O(f(n))$ with high probability (whp) (in $n$) to mean $O(cf(n))$ with probability at least $1-n^{-c}$ for $c\geq{1}$. in this stage; and (ii) each machine will hold $\Theta(\frac{n}{P})$ tasks whp at the end of this stage.
2.5. TD-Orch Evaluation

### v2

@@ 2.3 Scheduling Strategies
Direct Push. Another common strategy in distributed settings, such as in remote procedure calls, is to offload tasks to the machines that store the required data chunks. However, machines holding “hot” data chunks incur high communication cost (for receiving tasks) and high computation work (for executing received tasks), resulting in severe load imbalance. In the worst case, all $n$ tasks are pushed to the same chunk, resulting in $O(nD\sigma/\min\{D,P\})$ communication time and a competitive ratio of $O\left(\frac{P\sigma}{\min\{B,\sigma\}\min\{D,P\}}\right)$, which is also prohibitive.
@@
Theory-Guided Designs. Prior work on the Massively Parallel Computing (MPC) model [im2023massively] has proposed theoretically-efficient solutions to the orchestration problem (e.g., [im2023massively, goodrich2011sorting]). The procedure begins by sorting tasks according to the addresses of their required data chunks. Then, MPC broadcasting [ghaffari2019massively] is used to distribute each data chunk to the corresponding tasks. After task execution, a reverse broadcasting step writes the results back to the data chunks, followed by a reverse sorting step that restores tasks to their original order. It has been shown that the computation work is load balanced, and the communication time is $\Theta(\frac{n}{p}\log_{\frac{n}{P}}{P})$ [fish2015computational], which is optimal [roughgarden2018suffles]. Although these MPC schemes are asymptotically optimal, we show in §4 that their constant factors are non-negligible in practice, resulting in poor performance on real-world systems (e.g., [axtmann2015practical]) compared to empirical designs.
Empirical Designs. Numerous empirical designs (e.g., [vora2014aspire, li2020pegasus, gog2016firmament, murray2013naiad]), while not directly targeting task-data orchestration, employ practically efficient techniques in their scheduling, such as communication aggregation. However, few provide rigorous analysis of competitive ratios, and none offer tight bounds, limiting their ability to generalize to broader distributed systems and robustness to adversary workloads.
2.4 Prior Distributed Graph Processing
Classic distributed frameworks such as MapReduce [dean2008mapreduce] and Pregel [malewicz2010pregel, bu2014pregelix, han2015giraph] inspired many early distributed graph systems, including BLAS [bulucc2011combinatorial], GraphLab [low2012distributed], GraphMat [sundaram2015graphmat], Graphulo [gadepally2015graphulo], GraphX [gonzalez2014graphx], Mizan [khayyat2013mizan], Pegasus [kang2009pegasus], PowerGraph [gonzalez2012powergraph], PowerSwitch [xie2015sync], and Presto [venkataraman2013presto].
We classify more recent graph systems developed in the past decade into two major families: (i) Graph-algorithm-based: GeminiGraph [zhu2016gemini], Gluon [dathathri2018gluon], PowerLyra [chen2019powerlyra], FaasGraph [liu2024faasgraph], Phoenix [dathathri2019phoenix], ShenTu [lin2018shentu], and TopoX [li2019topox]; and (ii) Linear-algebra-based: Graphite [mofrad2020graphite], GraphPad [anderson2016graphpad], GraphScope [fan2021graphscope], GraphTap [mofrad2019efficient], and LA3 [ahmad2018la3].
Compared with prior graph-algorithm-based systems, our design, TDO-GP, achieves stronger provable and empirical load balance by using transit machines in its underlying TD-Orch to migrate contention. In contrast, prior mirror/ghost-node designs (e.g., [zhu2016gemini, gonzalez2012powergraph]) alleviate only limited contention and still behave largely as direct task-data exchanges.
Compared with prior linear-algebra-based designs, TDO-GP achieves stronger theoretical bounds and higher practical performance, by better exploiting graph sparsity than sparse-matrix approaches, by using work-efficient graph algorithms.
3 Main Design: TD-Orch
TD-Orch is a practically efficient orchestration framework with provable guarantees. Aligned with the original three-phase orchestration, TD-Orch refines Phase I (Read Data) by splitting it into two subphases—contention detection and task-data co-location. This yields a four-phase execution pipeline:
@@ 3 Main Design: TD-Orch
    Contention Detection: A preprocessing phase before the actual task-data co-location are carried out. The information of each task is sent to the machine storing its target data to gather the contention information for each data chunk. To preserve load balance, messages are passed along a communication forest network (§3.1), in the form of meta-tasks (§3.2).
@@
    Task-Data Co-location: Using the information gathered in Phase 1, tasks and data are redistributed across machines to achieve an optimized schedule that minimizes communication and computation time. (See §3.3.)
@@
    Data Write-backs: The updated value for any data modified by an executing task is written back to the original data location. (See §3.4.)
Key Design: Push-Pull. To address the challenges of direct push or pull, we must handle potential contention. Our method, TD-Orch, exploits the two directions of data flow: push, where a task context is sent to the machine(s) storing its required data chunks, and pull, where data chunks are fetched to the machine where the task resides. Such bidirectional strategies (referred to as distributed push-pull) are rarely emphasized in prior work. Only a few prior works demonstrate that this strategy can significantly improve load balance, but they are limited to PIM settings [kang2022pimtreevldb, kang2023pimtrie, zhao2025optimal, zhao2026pimzd, kang2025pim] and rely on an extremely powerful centralized coordinator (the host CPU). Our goal is to develop a more widely applicable push-pull technique that does not require a centralized coordinator.
Ideally, when contention information is available on any machine on the fly, distributed push-pull aggregates redundant information on one side (task/data) and dynamically chooses the flow direction toward the contended end. This dynamic decision should balance load with low communication.
3.1 Communication Forest for Handling Skew
Although the system assumes a fully connected network, tasks and data cannot be directly co-located, as this may cause severe load imbalance when hot data chunks exist. A circular challenge exists: we need a preprocessing Phase 1 to gather contention information for later load balancing, but Phase 1 itself must stay load-balanced before contention is even known.
In Phase 1 in TD-Orch, rather than exchanging all task-data in a single communication round, a tree-structured routing network in Fig. 2 is used to gradually push the information of tasks (which reside in the leaf machines and need to access a data chunk in the root machine of the tree) toward their target data across multiple rounds within Phase 1, moving up one tree level per round. This prevents tasks accessing the same hot data chunk from overloading a single machine.
@@ 3.1 Communication Forest for Handling Skew
Formally, a communication tree rooted at machine $i$ is a network with a balanced tree structure with $P$ leaves and fanout $F$, as shown in Fig. 2. Each of the $P$ leaf nodes represents a different physical machine, while internal nodes correspond to virtual transit machines, each of which is mapped to one of the $P$ machines via a hash function globally known to all machines. Notably, a physical machine can simultaneously serve as both a leaf and an internal node.
The communication tree is used to collect information about all tasks that need to access data chunks stored on Machine $i$. It enables efficient collection of task-related information (such as counting task occurrences) with load-balanced communication. The format of information passed along this network will be introduced in §3.2.
In the example in Fig. 2, Phase 1 needs three rounds (the height of trees). Suppose Machines 2 and 3 each have a task accessing data on Machine 1. In Round 1, each task sends a packed message (data address and task info) to its parents, VM(1,3)=PM3 and VM(1,4)=PM6. In Round 2, these messages are forwarded to VM(1,1)=PM5. In the final round, an aggregated message is delivered to Machine 1. This message aggregation prevents the overloading of a single machine.
A communication forest consists of $P$ such communication trees, each rooted at a distinct machine. All tasks send messages in parallel along the trees associated with the machines that store their target data chunks. Although the internal transit machines are statically chosen within the communication forest, each data chunk is stored at a random address (§2.2). This randomization in the corresponding tree selection is equivalent to dynamically selecting transit nodes in both theoretical analysis and practical implementation.
We use $F=\Theta\left(\log P/\log\log P\right)$ in TD-Orch based on the theoretical analysis in §3.5, which also yields strong practical performance as shown in §4and §6.
Figure 2: A communication forest with fanout $F=2$ for $P=8$ machines. Green nodes are source machines (task storage), red nodes are destination machines (data storage), and blue nodes are intermediate transit virtual machines (VMs) with IDs $(\text{root},\text{BFS\_id})$, mapped to physical machines (PMs) via example hash function $h(x,y)=(x+3y)\bmod 8+1$. Dashed edges mean that PM does not need to send messages to itself.
3.2 Meta-Task Structure
This section describes the messages passed through the communication forest during Phase 1. These messages are used to support co-location in Phase 2 and should include (i) the reference count for each data chunk, and (ii) the locations of the tasks requesting them. They should also be aggregatable to prevent unbounded message size growth at any machine.
To meet these goals, we introduce a meta-task structure—parameterized by hyperparameter $C$ and attribute LL—to serve as the message payload within the tree-structured network. At a high level, if the reference count of a data chunk is $\leq C$, the meta-task contains the full contexts of all requesting tasks. If the reference count exceeds $C$, the meta-task includes only aggregated metadata, including the count and the locations of tasks, thereby preventing unbounded message growth.
The attribute $L$, which serves as a level indicator, enables hierarchical aggregation of meta-tasks. Raw task contexts are treated as $L_{0}$ meta-tasks. During message passing in the communication forest, if an intermediate machine receives more than $C$ $L_{i}$ meta-tasks, it locally stores them and creates a new $L_{i+1}$ meta-task. This new $L_{i+1}$ meta-task contains two pieces of information: (i) the aggregated reference count, and (ii) the addresses of the machines where the $C$ underlying $L_{i}$ meta-tasks are stored. Only this aggregated $L_{i+1}$ meta-task will be forwarded to its parent machine.
@@ 3.2 Meta-Task Structure
A meta-task set, as shown in Figure˜3, is a set structure containing multiple meta-tasks. The number of meta-tasks in a meta-task set with the same $L$ value should not exceed $C$.
@@
A merge operation of two meta-task sets is proposed to support message aggregation. A new meta-task set is generated, containing all meta-tasks from both of the original sets. If the number of meta-tasks at a certain level $L_{i}$ in the new set exceeds $C$, all $L_{i}$-meta-tasks will be popped out of the new set and stored locally on the current machine, and then a new $L_{l+1}$-meta-task will be pushed into the set. The merge will be executed from the lowest level up to the highest level, until the number of meta-tasks at each level is at most $C$.
Example Fig. 4 with $C=3$ shows the merge process. First, $L_{0}$-meta-tasks are merged; since their count exceeds $C$, they are locally stored and an $L_{1,\text{new}}$ is created. Next, the (now >C>C) $L_{1}$-meta-tasks, including $L_{1,\text{new}}$, are similarly stored and produce an $L_{2,\text{new}}$. There are $\leq C$ $L_{2}$-meta-tasks, so all are emitted to the output set. Continue this until the highest level.
This merging process guarantees that each meta-task set has bounded size $C\log_{C}n$, while containing sufficient location information to enable efficient task-data exchange in Phase 2.
We use $C=\Theta(B/\sigma)$ in TD-Orch following the theory-guided design in §3.5, which also leads to strong practical performance as shown in §4 and §6.
Figure 3: A meta-task set, zooming in $L_{0}$ and $L_{i}$ ($i\geq 1$) meta-tasks. Deeper colors indicate higher levels. $L_{0}$ contains original task contexts, while $L_{i}$($i\geq{1}$) contains a pointer to a $L_{i-1}$-meta-task array stored on a machine.
@@
Figure 4: Merging two meta-task sets with $C=3$. Meta-tasks on lowest level $L_{0}$ are merged first and generate $L_{1,new}$. Then $L_{1}$-meta-tasks (including both the received $L_{1}$ and the newly generated $L_{1,new}$) are merged and generate $L_{2,new}$. Continue the process until the top level of $L_{4}$.
3.3 Distributed Push-Pull for Co-location
After Phase 1, each data chunk receives a meta-task set that contains all information needed for task-data co-location. In Phase 2, we introduce distributed push-pull, a strategy that exploits the bidirectional flow of tasks and data to resolve potential contentions on data chunks.
Push. If the reference count of a data chunk is $\leq{C}$, then all task contexts requesting this chunk have already been delivered as $L_{0}$ meta-tasks to the machine storing the chunk. Thus, for these uncontended tasks, the push process—sending tasks to data—is already completed during Phase 1.
Pull. For contended tasks, where the reference count of a data chunk exceeds $C$, tasks are stored on intermediate transit machines in the communication forest. A pull process—sending data (copies) to the tasks—is required in Phase 2. The location information in meta-task sets is used to guide the pull.
As an example, suppose a data chunk receives a meta-task set with the highest level $L_{i}$. It first broadcasts a copy of the chunk to all $\Theta(C)$ machines storing the $L_{i-1}$-meta-tasks, then chases pointers and broadcasts the data to the $\Theta(C^{2})$ machines storing the $L_{i-2}$-meta-tasks, continuing until all $L_{0}$ meta-tasks (the original task contexts) have received a copy.
This broadcasting follows a path of an inherent tree-structured network formed by a meta-task set, which we define as the meta-task tree. This tree is more compact than the communication tree, forming a full $C$-ary tree structure with all leaf machines containing $L_{0}$-meta-tasks and no empty leaves.
Key Takeaway. Bi-directional distributed push-pull is a key design that enables load-balanced execution under contention while maintaining efficient communication overheads. It combines push and pull: uncontended tasks are directly pushed to their target data, whereas contended tasks pull their data gradually, level by level, along the inherent meta-task tree.
3.4 Write Backs
If tasks require write-backs that are merge-able (defined below), they can be efficiently aggregated, allowing the write-backs from multiple tasks to the same data chunk to be combined. In Phase 4, these write-backs are propagated along the reverse path of the meta-task tree defined in §3.3, applying this aggregation process during message passing.
@@ 3.4 Write Backs
3.5 Formal Guarantees
We use the BSP model [valiant1990bridging, valiant1990general, valiant1988optimally] to provide theoretical guarantees for our design. Readers less interested in the theory can directly refer to the conclusions in Key Takeaway.
@@ 3.5 Formal Guarantees
Our main Theorem˜1 proves TD-Orch to be asymptotically optimal compared with the lower bounds proved in [roughgarden2018suffles]. Please refer to Appendix B for its proof.
Theorem 1 (Main Theorem).
@@
Assume $n=\Omega(\frac{PB\log^{3}{P}}{\log\log{P}})$ and $n=O(P^{\epsilon})$ for some $\epsilon>1$. If each machine holds $\Theta(n/P)$ tasks when a stage begins, and write backs are merge-able, then (i) TD-Orch takes $O(\frac{nD}{P}(\min\{B,\sigma\}+\log_{\frac{n}{P}}{P}))$ communication time whp  [^2] We use $O(f(n))$ with high probability (whp) (in $n$) to mean $O(cf(n))$ with probability at least $1-n^{-c}$ for $c\geq{1}$. in this stage; and (ii) each machine will execute $\Theta(\frac{n}{P})$ tasks whp and hold them at the end of this stage.
Key Takeaway. (i) Optimal Communication: TD-Orch incurs optimally minimal communication time as $P$ scales up. (ii) Optimal Computation: TD-Orch balances computation works, maximizing system throughput. (iii) Inductive Execution: Task storage remains load-balanced after each stage, enabling efficient execution of subsequent stages.
3.6 Comparison Against Prior Methods
Against Direct Push/Pull. TD-Orch achieves the optimal theoretical bounds as shown in §3.5, outperforming load-balance-oblivious methods.
Against Prior Theory-Guided Methods. Though asymptotically the same (both optimal), TD-Orch requires only $2$ sweeps over the communication forest, whereas MPC sorting [im2023massively, goodrich2011sorting] incurs $\geq{3}$ passes. This reduction reflects the potential of TD-Orch for strong practical efficiency.
Against Prior Empirical Designs. TD-Orch (i) provides provable guarantees, and (ii) explicitly targets the orchestration abstraction, rather than a fused design of scheduling inside specific end-to-end applications.
4 Case Study I: Key-Value Store
A natural fit for our orchestration framework is a key-value store, which we implement atop TD-Orch using a concurrent distributed hash table as the underlying storage structure.